OpenStudy (anonymous):

On your birthday, you deposit $540.00 in an account that pays 6% interest, compounded annually. How much is in the account 3 years later? (1 point) a.$637.20 b.$543.18 c.$643.15 d.$1,717.20

OpenStudy (anonymous):

remember the formula for compount interest here we have principal P =540 interest rate r= 0.06 compounded anually so n=1 and time is t=3

OpenStudy (anonymous):

Compound interest formula is M = P(1 + i)^n Where M = the final amount P = the principal i = interest rat per year n = number of years invested M = 540(1 + 0.06)^3

OpenStudy (anonymous):

You first do what's in the parenthesis. so 1 + .06 =?

OpenStudy (anonymous):

that is a decimal before the 06

OpenStudy (anonymous):

okay, now you calculate 1.06 to the power of 3

OpenStudy (anonymous):

powers are the number multiplied by itself, that many times so it would be the same as saying (1.06)(1.06)(1.06)

OpenStudy (anonymous):

1.06 * 1.06 =?

OpenStudy (anonymous):

you're adding them together and you need to be multiplying them

OpenStudy (anonymous):

very good, now multiply that number by the amount being invested which is 540

OpenStudy (anonymous):

and since you're dealing with money you'll need to round to the nearest 2 decimal places.

OpenStudy (anonymous):

so that number rounded off should give you?

OpenStudy (anonymous):

You would be correct.

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The simple interest formula is I = Prt Where I = the interest P = the principal r = the interest rate t = the time it takes you to pay of the loan I = (500)(0.03)(2)

OpenStudy (anonymous):

Solve the problem as it sits to give you A) Add the amount borrowed with the answer for A) to give you B)

OpenStudy (anonymous):

the parenthesis mean those numbers are being multiplied

OpenStudy (anonymous):

you need to multiple one more number into that

OpenStudy (anonymous):

that's your A)

OpenStudy (anonymous):

$30 in interest will be paid

OpenStudy (anonymous):

B) is the sum of the principal and the interest

OpenStudy (anonymous):

Unless you have another question about these problems I'd say were done. I hope that helped